3 apples cost as much as 4 pears, 3 pears cost as much as 2 oranges.
How many apples cost the same as 72 oranges

Answer by Earlsdon(6103)

Let A = cost of 1 apple.
Let P = cost of 1 pear.
Let (Or) = cost of 1 orange.
(1) 3A = 4P So, P = (3/4)A Substitute into (2) below.
(2) 3P = 2(Or)
(2a) 3(3/4)A = 2(Or) Divide both sides by 2.
(Or) = (9/8)A Multiply both sides by 72.
72(Or) = 72(9/8)A
72(Or) = 81A
So, 81 Apples cost the same as 72 Oranges.